Michael Jones, Concord & Jones Fabricators' production manager, has just received the company's sales budget for the first quarter.

January

February

March

April

May

June

21,000

29,000

32,000

33,000

22,000

34,000


Company policy requires an ending finished goods inventory each month that will meet 30% of the following month's sales volume. Michael plans to have 2,400 finished bricks at a cost of $49,280 in inventory at the beginning of the year.

Prepare Concord & Jones's production budget for the first quarter.

January

February

March

Quarter

select an opening production budget item                                                          Budgeted ending inventoryTotal units requiredBeginning inventoryBudgeted unit salesBudgeted production

enter a number of units enter a number of units enter a number of units enter a number of units

select a production budget item                                                          Budgeted productionTotal units requiredBudgeted ending inventoryBudgeted unit salesBeginning inventory

enter a number of units enter a number of units enter a number of units enter a number of units

select a summarizing line for the first part                                                          Beginning inventoryBudgeted ending inventoryTotal units requiredBudgeted productionBudgeted unit sales

enter a total number of units for the first part enter a total number of units for the first part enter a total number of units for the first part enter a total number of units for the first part

select a production budget item                                                          Beginning inventoryBudgeted unit salesBudgeted ending inventoryTotal units requiredBudgeted production

enter a number of units enter a number of units enter a number of units enter a number of units

select a closing production budget item                                                          Budgeted ending inventoryBeginning inventoryTotal units requiredBudgeted productionBudgeted unit sales

enter a total number of units enter a total number of units enter a total number of units enter a total number of units

eTextbook and Media
Answer #1
January February March Quarter
Budgeted unit sales                              21,000                              29,000                              32,000                              82,000
Add: Budgeted ending inventory 29,000*30% = 8,700 32,000*30% = 9,600 33,000*30% = 9,900 33,000*30% = 9,900
Total units required                              29,700                              38,600                              41,900                              91,900
Less: Beginning inventory                                2,400 29,000*30% = 8,700 32,000*30% = 9,600                                2,400
Budgeted production                              27,300                              29,900                              32,300                              89,500

You can reach me over comment box, if you have any doubts. please rate this answer